# Theoretical and Empirical Insights into the Origins of Degree Bias in Graph Neural Networks

### v1

@@ 3.3.1 Symmetric Graph Filter
3.3.2 Random Walk Graph Filter
@@
We now demonstrate that during each step of training $\overline{\text{RW}}$ with gradient descent, compared to $\overline{\text{SYM}}$, the loss of low-degree nodes in $S$ is not necessarily adjusted more slowly. We define $\forall l\in[L],\chi_{i}^{(l)}\in\mathbb{R}^{|B[t]|}$, where for $m\in B[t],\left(\chi_{i}^{(l)}[t]\right)_{m}=\mathbb{E}_{j\sim{\cal N}^{(l)}(i),k\sim{\cal N}^{(l)}(m)}\left[{\bm{X}}_{j}{\bm{X}}_{k}^{T}\right]$. In effect, $\left(\chi_{i}^{(l)}[t]\right)_{m}$ captures the expected similarity between the raw features of nodes $j$ and $k$ with respect to the $l$-hop random walk distributions of $i\in{\cal V}$ and $m\in B[t]$.
@@
$\left|\ell[t+1](\overline{\text{RW}}|i,c)-\ell[t](\overline{\text{RW}}|i,c)\right|$ $\leq\sqrt{2}\eta\left\|\epsilon[t]\right\|_{F}\sum_{l=0}^{L}\left\|\chi_{i}^{(l)}[t]\right\|_{2}.$
For $\overline{\text{RW}}$, the change (either increase or decrease) in loss for $i$ after an arbitrary training step does not necessarily have a smaller magnitude if $i$ is low-degree. However, the $L$-hop neighborhoods of low-degree nodes still often have less overlap with the neighborhoods of training nodes, which can constrain the rate at which the loss for $i$ changes. We confirm these findings empirically in Figure 2 and §F. For all the datasets, the blue curve for RW has a less steep rate of decrease than the blue curve for SYM as the number of epochs increases. However, for RW itself, the orange curve generally descends more quickly than the blue curve, with the exception of the heterophilic chameleon and squirrel datasets, for which the features of nodes in the neighborhoods of high-degree nodes and training nodes are dissimilar. These findings support hypothesis (H2). Our results for $\overline{\text{RW}}$ may also apply to ATT when low-degree nodes are generally attended to less.
3.4 Achieving Maximum Training Accuracy
@@
Per [51], because the expressive power of SYM, RW, and ATT are limited by the WL test, an upper bound on their training accuracy is the accuracy of a majority voting-classifier $\textsc{MAJ}_{\text{WL}}$ applied to WL node colorings (for details on how to compute colorings, cf. §II.A and §VI.B of [61]). In particular, if the WL test produces the colors ${\bm{c}}\in\mathbb{K}^{|S|}$ for nodes in $S$, $\textsc{MAJ}_{\text{WL}}$ predicts node $i$ to have the label $\widehat{{\bm{Y}}}_{i}=\textsc{MAJ}_{\text{WL}}(i,{\bm{X}},{\bm{A}})=\text{mode}\{{\bm{Y}}_{j}|j\in{\cal V},{\bm{c}}_{j}={\bm{c}}_{i}\}$. However, Figure 4 and §H reveal that as the number of training epochs increases, the training accuracy of SYM, RW, and ATT reach the accuracy of $\textsc{MAJ}_{\text{WL}}$ Because the accuracy of $\textsc{MAJ}_{\text{WL}}$ is often close to 100%, our experiments suggest that insufficient expressive power likely does not contribute to degree bias, drawing doubt to hypothesis (H7).
@@
$\frac{\partial\ell[t]}{\partial{\bm{W}}^{(l)}[t]}(B[t])$ $={\bm{X}}^{T}\left({\bm{P}}^{l}_{\text{rw}}[t]\right)^{T}\epsilon[t].$
@@
4 Principled Roadmap to Address Degree Bias
Future work can build on our theoretical and empirical insights to propose alleviation strategies for degree bias. In particular, our findings reveal that any strategies should target the following criteria:
  • Maximizing the inverse collision probability of low-degree nodes (e.g., via edge augmentation for low-degree nodes).
  • Increasing the $L$-hop prediction homogeneity of low-degree nodes (e.g., by ensuring similar label densities in the neighborhoods of low and high-degree nodes).
  • Minimizing distributional differences (e.g., differences in expectation, variance) in the representations of low and high-degree nodes.
  • Reducing training discrepancies with regards to the rate at which GNNs learn for low vs. high-degree nodes.
5 Conclusion
Our theoretical analysis aims to unify and distill previously-proposed hypotheses for the origins of GNN degree bias. We prove that high-degree test nodes tend to have a lower probability of misclassification and that degree bias arises from a variety of factors associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors). Furthermore, we show that during training, some GNNs may adjust their loss on low-degree nodes more slowly; however, GNNs often achieve their maximum possible training accuracy and are trivially limited by their expressive power. We validate our theoretical findings on 8 real-world networks. Finally, based on our theoretical and empirical insights, we describe a roadmap to alleviate degree bias. More broadly, we encourage research efforts that unveil forms of inequality reinforced by GNNs.
Acknowledgments and Disclosure of Funding
This work was supported by NSF 2303037 and NSF 1829071.
Limitations
@@ 5 Conclusion
Our theoretical analysis is limited to linearized message-passing GNNs. While this is a common practice in the literature [45, 6, 34], it is a strong simplifying assumption. We empirically validate our theoretical findings on GNNs with non-linear activation functions, but our paper does not address possible sources of degree bias related to these non-linearities. Our assumptions that GNNs generalize in expectation (Theorem 1) and the variance of node representations is finite (Theorems 2 and 3) are not overly strong assumptions in practice.
@@ 14. Answer: [Yes]
  15. Justification: Details to reproduce experiments are provided in the main body of the paper, §D, and §C. The code is released on GitHub.
@@ 18. Answer: [Yes]
  19. Justification: The code is released on GitHub. The data are available through PyTorch Geometric [14], and data processing details are provided in §C.
@@ 44. Answer: [Yes]
  45. Justification: See §C. The original paper that produced each dataset is cited. Links to libraries and datasets are included; the versions of software dependencies can be found in the requirements.txt file on GitHub. We also provide the name and a link to the license (which includes terms of use and copyright information, where applicable) of each library and dataset.
@@ 48. Answer: [Yes]
  49. Justification: We release our code on GitHub under an MIT license. We provide a README that explains how to reproduce our experiments. Training, license compliance, and consent details are included in §D. Limitations are discussed in §Limitations.

### v2

@@ Appendix G Additional Inverse Collision Probability Plots
Appendix H Training-Time Degree Bias: Random Walk Graph Filter
@@
We now demonstrate that during each step of training $\overline{\text{RW}}$ with gradient descent, compared to $\overline{\text{SYM}}$, the loss of low-degree nodes in $S$ is not necessarily adjusted more slowly. We define $\forall l\in\mathbb{N}_{\leq L},\chi_{i}^{(l)}\in\mathbb{R}^{|B[t]|}$, where for $m\in B[t],\left(\chi_{i}^{(l)}[t]\right)_{m}=\mathbb{E}_{j\sim{\cal N}^{(l)}(i),k\sim{\cal N}^{(l)}(m)}\left[{\bm{X}}_{j}{\bm{X}}_{k}^{T}\right]$. In effect, $\left(\chi_{i}^{(l)}[t]\right)_{m}$ captures the expected similarity between the raw features of nodes $j$ and $k$ with respect to the $l$-hop random walk distributions of $i\in{\cal V}$ and $m\in B[t]$.
@@
$\left|\ell[t+1](\overline{\text{RW}}|i,c)-\ell[t](\overline{\text{RW}}|i,c)\right|$ $\leq\sqrt{2}\eta\left\|\epsilon[t]\right\|_{F}\sum_{l=0}^{L}\left\|\chi_{i}^{(l)}[t]\right\|_{2}.$ (66)
For $\overline{\text{RW}}$, the change (either increase or decrease) in loss for $i$ after an arbitrary training step does not necessarily have a smaller magnitude if $i$ is low-degree. However, the $L$-hop neighborhoods of low-degree nodes still often have less overlap with the neighborhoods of training nodes, which can constrain the rate at which the loss for $i$ changes. We confirm these findings empirically in Figure 2 and §F. For all the datasets, the blue curve for RW has a less steep rate of decrease than the blue curve for SYM as the number of epochs increases. However, for RW itself, the orange curve generally descends more quickly than the blue curve, with the exception of the heterophilic chameleon and squirrel datasets, for which the features of nodes in the neighborhoods of high-degree nodes and training nodes are dissimilar. Therefore, models do not learn more rapidly for high-degree nodes under heterophily. These findings support hypothesis (H2). Our results for $\overline{\text{RW}}$ may also apply to ATT when low-degree nodes are generally attended to less.
Appendix I Achieving Maximum Training Accuracy
@@
Per [57], because the expressive power of SYM, RW, and ATT are limited by the WL test, an upper bound on their training accuracy is the accuracy of a majority voting-classifier $\textsc{MAJ}_{\text{WL}}$ applied to WL node colorings (for details on how to compute colorings, see §II.A and §VI.B of [68]). In particular, if the WL test produces the colors ${\bm{c}}\in\mathbb{K}^{|S|}$ for nodes in $S$, $\textsc{MAJ}_{\text{WL}}$ predicts node $i$ to have the label $\widehat{{\bm{Y}}}_{i}=\textsc{MAJ}_{\text{WL}}(i,{\bm{X}},{\bm{A}})=\text{mode}\{{\bm{Y}}_{j}|j\in{\cal V},{\bm{c}}_{j}={\bm{c}}_{i}\}$. However, Figure 15 and §J reveal that as the number of training epochs increases, the training accuracy of SYM, RW, and ATT reach the accuracy of $\textsc{MAJ}_{\text{WL}}$. Because the accuracy of $\textsc{MAJ}_{\text{WL}}$ is often close to 100%, our experiments suggest that insufficient expressive power likely does not contribute to degree bias, drawing doubt to hypothesis (H7).
@@
$\frac{\partial\ell[t]}{\partial{\bm{W}}^{(l)}[t]}(B[t])$ $={\bm{X}}^{T}\left({\bm{P}}^{l}_{\text{rw}}[t]\right)^{T}\epsilon[t].$ (67)
@@
Appendix J Additional Training Loss Plots
Appendix K Connecting Inverse Collision Probability to Node Degree
Our theoretical analysis may be improved upon by establishing a rigorous connection, or a lack thereof, between the inverse collision probability of a node and its degree.
Via some preliminary analysis, we find that it is possible to express the inverse collision probability of $i$ (i.e., equivalently express the sum of $l$-hop collision probabilities) in terms of ${\bm{D}}_{ii}$. In particular, we can show: $\sum_{l=0}^{L}\alpha_{i}^{(l)}=\alpha_{i}^{(0)}+\sum_{l=1}^{L}\sum_{j\in{\cal V}}[({\bm{P}}^{l}_{rw})_{ij}]^{2}=1+\frac{1}{D^{2}_{ii}}\sum_{l=1}^{L}\sum_{j\in{\cal V}}[\sum_{k\in{\cal N}(i)}({\bm{P}}^{l-1}_{rw})_{kj}]^{2}$. As before, we can see that the inverse collision probability is larger (and thus $R_{i,c^{\prime}}$ is larger) when ${\bm{D}}_{ii}$ is larger and $\sum_{l=1}^{L}\sum_{j\in{\cal V}}[\sum_{k\in{\cal N}(i)}({\bm{P}}^{l-1}_{rw})_{kj}]^{2}\in o({\bm{D}}_{ii}^{2})$. However, because $\sum_{k\in{\cal N}(i)}$ depends on ${\bm{D}}_{ii}$, this expression does not completely isolate the impact of ${\bm{D}}_{ii}$ on the inverse collision probability. A similar expression can be derived for SYM by expressing: ${\bm{P}}_{sym}={\bm{D}}^{\frac{1}{2}}{\bm{P}}_{rw}{\bm{D}}^{-\frac{1}{2}}$.
Alternatively, one may consider $\sum_{l=0}^{L}\alpha_{i}^{(l)}\leq\sum_{l=0}^{\infty}\alpha_{i}^{(l)}$, and then plug in the steady-state probabilities of uniform random walks on graphs. However, as $l\to\infty$, $({\bm{P}}^{l}_{rw})_{ij}=\frac{{\bm{D}}_{jj}}{2|{\cal E}|}$ only depends on ${\bm{D}}_{jj}$ (not ${\bm{D}}_{ii}$, as desired). It is also possible to upper bound the inverse collision probability as: $1/\sum_{l=0}\alpha_{i}^{(l)}\leq 1/(\alpha_{i}^{(0)}+\alpha_{i}^{(1)})=1/(1/{\bm{D}}_{ii}+1)$, which is in terms of ${\bm{D}}_{ii}$; however, we cannot use this upper bound to lower bound $R_{i,c^{\prime}}$. The collision probabilities themselves are intimately related to more global properties of graphs; for instance, via eigendecomposition, $[({\bm{P}}^{l}_{rw})_{ij}]^{2}\leq\lambda^{2l}$, where $\lambda<1$ is the eigenvalue of ${\bm{P}}_{rw}$ with the largest magnitude [36]. Then, the inverse collision probability is strictly greater than $\frac{1}{\sum_{l=0}^{L}|{\cal V}|\lambda^{2l}}$.
This being said, our paper argues that the inverse collision probability is a more fundamental quantity that influences what the fair graph learning community has termed “degree bias” than degree alone, which we have shown is positively associated with the inverse collision probability.
Appendix L Limitations and Future Directions
@@
Our theoretical analysis is limited to linearized message-passing GNNs. While this is a common practice in the literature [51, 7, 39], it is a strong simplifying assumption. We empirically validate our theoretical findings on GNNs with non-linear activation functions, but our paper does not address possible sources of degree bias related to these non-linearities, which would be interesting to investigate in future work. Towards this, a possible option is to assume that node features are drawn from a Gaussian distribution and derive precise high-dimensional asymptotics for degree bias in non-linear GNNs using the Gaussian equivalence theorem, as in [1]. Our assumptions that GNNs generalize in expectation (Theorem 1) and the variance of node representations is finite (Theorems 2 and 3) are not overly strong assumptions in practice.
Furthermore, our paper focuses on node classification. However, our findings readily lend insight into the origins of degree bias in link prediction. For example, if one uses node representations and an inner-product decoder to predict links between nodes, our results indicate that:
  • In the random walk filter case, link prediction scores between low-degree nodes will suffer from higher variance because low-degree node representations have higher variance (Theorem 2). Hence, Theorem 1 suggests that predictions for links between low-degree nodes will have a higher misclassification error.
  • In the symmetric filter case, our proof of Theorem 3 suggests that the link prediction scores between high-degree nodes will be over-calibrated (i.e., disproportionately large) because high-degree node representations have a larger magnitude (i.e., approximately proportional to the square root of their degree). Hence, over-optimistic and possibly inaccurate links will be predicted between high-degree nodes.
The labels and evaluation for link prediction can confound intuition. Unlike node classification, the labels for link prediction (i.e., the existence or not of a link) make the task naturally imbalanced with respect to node degree; high-degree nodes have a much higher rate of positive links than low-degree nodes. This association between degree and positive labels can influence the misclassification error. Ultimately, more rigorous theoretical analysis and experimentation are needed to confirm the hypothesized implications of node degree for link prediction performance. Similarly, more research is required to understand the implications of our findings for degree bias in the context of graph classification.
Furthermore, our theoretical analysis does not encompass heterogeneous graphs. In our literature survey, we cover works that establish the issue of degree bias for knowledge graph predictions and embeddings (e.g., [5, 43]). Our theoretical analysis is general and covers diverse message-passing GNNs, and can be extended to heterogeneous networks if messages aggregated from different edge types are subsequently linearly combined. In this setting, $R_{i,c^{\prime}}$ can be computed as the sum of the prediction homogeneity quantities $(\sum_{l=0}^{L}\beta_{i,c^{\prime}}^{(l)})^{2}$ for each edge type divided by the sum of the collision probability quantities $\sum_{j\in{\cal V}}[({\bm{P}}_{rw}^{l})_{ij}]^{2}$ for each edge type.
@@ 14. Answer: [Yes]
  15. Justification: Details to reproduce experiments are provided in the main body of the paper, §D, and §C. The code is released via Github.
@@ 18. Answer: [Yes]
  19. Justification: The code is released via Github. The data are available through PyTorch Geometric [15], and data processing details are provided in §C.
@@ 44. Answer: [Yes]
  45. Justification: See §C. The original paper that produced each dataset is cited. Links to libraries and datasets are included; the versions of software dependencies can be found in the requirements.txt file via Github. We also provide the name and a link to the license (which includes terms of use and copyright information, where applicable) of each library and dataset.
@@ 48. Answer: [Yes]
  49. Justification: We release our code via Github under an MIT license. We provide a README that explains how to reproduce our experiments. Training, license compliance, and consent details are included in §D. Limitations are discussed in §L.